# Improving Bilinear RNNs with Closed-loop Control

### v4

@@ 3.1 Forcing Forgetting for Long-range Modeling
Extensive work [61, 92, 76] has shown that positional encoding is crucial for language models to generalize from short pretraining contexts to unconstrained inference lengths. While gated linear recurrent structures themselves can be viewed as the continuous accumulation of bias [92], enabling these models to extrapolate effectively by learning how to forget. Recent studies [20, 110, 9] suggest that pretraining lengths (e.g., 2K or 4K tokens) are insufficient to fill the model’s state capacity, thus making the forget gate close to 1, forcing the model to learn how to forget. As shown in Table LABEL:tab:forcing_forget, we evaluate representative gating initialization methods. For models like GLA [107], the forget gate is initialized to $(0,1)$ and retains all incremental memories. For MetaLA [22], HGRN2 [79], and GSA [112], the sum of the forget and input gates is constrained to 1, achieving a relative balance in system memory capacity. Experimental results in §4.2 indicate that breaking this balance is necessary. To this end, Comba follows Mamba2’s design by constraining the forget gate $\alpha_{t}$} to close to 1, while separately initializing the input gate $\beta_{t}$} to $(0,1)$ [108]. Additionally, we set the strength of state feedback correction is $\tilde{\beta}_{t}={b}\odot\beta_{t}$, where b{b} is computed by $\operatorname{Sigmoid}$ function to be constrained in interval $(0,1)$ to ensure that the feedback strength is weaker than the incremental information.
@@ 3.2 Comba with Chunk-wise Parallel
Based on Eq. 5, Comba can be implemented recursively, enabling constant memory usage and $\mathcal{O}(1)$ time complexity during the inference stage, where the Python-style pseudocode is provided in App. B.1. However, the naive recursive implementation in PyTorch lacks sufficient matrix multiplication and thread parallelism [25], resulting in unacceptable overhead to pretraining. Referring to DeltaNet [105], we optimize Comba through chunk-wise parallelism. App. B.3 presents an alternative form to fuse feedback decay factor $b$ into $\bm{k}$ as $\bm{p}=b\bm{k}$ for a flexible invocation.
@@
Eq. 6 involves matrix-matrix products at each time step, i.e., $\bm{S}_{[t]}^{0}@\bm{D}_{[t]}^{r}$, preventing parallelization in the sequence level. Then, we employ the WY representation [13] to eliminate these terms:
@@
To maximize hardware efficiency, we apply the UT transform [51] to Eq. 7-8 to reduce non-matmul FLOPs, which is crucial to enable better hardware utilization during training:
@@
The inverse of a lower triangular matrix can be efficiently computed through an iterative row-wise approach by forward substitution in Gaussian elimination [37] and maintain data in $\operatorname{float32}$. Notably, Comba computes the inverse matrix once in Eq. 10 (twice in the original Gated-DeltaNet). This is mainly attributed to the introduction of a new form of mathematical induction in deriving the WY representation in Eq. 7, leading to a more concise formulation and resulting in speedup [^3] During the forward pass, Comba alleviates the main bottleneck of inverse matrix computation with $\mathcal{O}(d^{3})$ in the original Gated-DeltaNet, yielding a 40% speedup. In the backward pass, performance gains diminish due to the reuse of cached inverse matrices $\bm{M}$. However, in large-scale models where recomputation is required, Comba is expected to offer significant performance advantages. Moreover, this structured modification applies to almost all models with Delta rule. Recently, inspired by Comba, Gated-DeltaNet has also been upgraded to a single inversion in flash-linear-attention, resulting in a significant speedup..
@@ 3.3 Neural Architecture
As shown in Fig. 2, we present the architecture for Comba families, where $\{a,b,c,d\}$ are trainable scalars (experimental results indicate that data dependency is not required). Following prior work [22, 105, 38, 32], we introduce short convolutions to $\bm{qkv}$ to incorporate token shift to improve the model’s retrieval capacity. We also retain feature map operations [98] and utilize the $\operatorname{SiLU}$ function to approximate the exponential kernel in the softmax attention. To further stabilize training, we apply L2 normalization to $\bm{qk}$ and employ a $\operatorname{Sigmoid}$-based gating mechanism. Additionally, we explore a hybrid architecture [89, 105, 80] by integrating Comba layers directly with softmax attention.
@@ 4.1 Operator Efficiency Analysis
Figure 3: Operator speed evaluated on the Triton-Testing-Benchmark [93] (fwd and bwd) in single A800-80G GPU.
As shown in Fig. 3, we compared the speeds of various operators in both forward and backward processes. The recurrent Comba in PyTorch [71] incurs significant computational overhead, limiting its scalability for large-scale pretraining. Flash-attention [25] achieves the fastest speed for shorter sequences (e.g., 1024), but its quadratic complexity results in decreasing efficiency as sequence length increases. Among four modern RNN operators, Comba shows nearly 40% speed improvement in the forward process over Gated-DeltaNet due to a more efficient formula structure. GLA suffers from slower operator speed due to the use of diagonal gating matrices. Although RetNet achieves the fastest speed, it falls behind other models in performance due to the lack of data-dependent memory management (Table 6). Overall, Comba shows potential as a foundational framework.
@@ 4.2 Language Modeling & Architecture Ablation
Table 6: Zero-shot performance of 340M and 1.3B models trained on SlimPajama [85] datasets. The commonsense Reasoning task is evaluated by lm-evaluation-harness [33] and the recall-intensive task follows prefix-linear-attention [3] with 2K input tokens. * Some of the baseline results are from [108] and [29].
@@
As shown in the left half of Table 6, (i) most recursive models outperform transformers in commonsense reasoning tasks, due to their recursive structure that resembles a chain of thought [100]. (ii) The SPLR structure outperforms both the IPLR and DPLR structures in two model sizes and achieves the highest computational efficiency. (iii) Output correction, i.e., $\left<q,dk\right$>, significantly reduces perplexity, enhancing memory utilization during question answering and improving performance across various metrics. (iv) We find that the Mamba architecture design is suboptimal. MLP, as a special (non-Hebbian) key-value memory [35, 34], complements the key-value associative memory in the state, which is particularly important for tasks such as inference. (v) Although Mamba’s multi-value attention model aids efficient key-value memory storage [59], it sacrifices performance compared to standard multi-head attention. (vi) We discover that the initialization of $d$ should be chosen differently for model scales, e.g., $d=0.02$ for 340M models and $d=1$ for 1.3B models, as smaller models are more prone to incorrect gradient descent directions.
@@
Inspired by previous hybrid architectures [105, 89, 14, 28, 60, 19], we empirically replace the second-to-last layer in every eight layers of Comba with softmax attention to boost the model’s recall ability. In Table 7, (i) we find that nearly all hybrid architectures outperform transformers even with a few quadratic layers. (ii) The best overall performance comes from the GSA-H [112], as GSA itself is a type of intra-layer hybrid architecture. Computationally, it can be expressed as two GLA operations followed by a softmax operation to address attention sparsity issues [92], thus improving recall ability. In the future, we plan to combine GSA with Comba for a flexible hybrid.
@@
Table 8: Performance on LongBench [4] tasks with 10K length based on lm-evaluation-harness [33].
@@ 4.3 Vision Modeling
Table 9: Performance on the ImageNet-1K [27] classification, compared to Vision Mamba [115] (linear), DeiT [94] (quadratic), and Agent Attention [44] (sparse). Model Res. Params. FLOPs Top-1 DeiT-T 224*224 5.7(MB) 1.2(G) 72.2% Agent-T 224*224 6.0 1.2 74.9(+2.7) Vim-T 224*224 7.0 1.5 76.1(+3.9) Comba-T 224*224 5.8 1.1 76.3(+4.1) DeiT-S 224*224 22.1 4.6 79.8 Agent-S 224*224 22.7 4.6 80.5(+0.7) Vim-S 224*224 26.0 5.1 80.3(+0.5) Comba-S 224*224 22.6 4.4 80.5(+0.7)    Table 10: Performance on the object tracking datasets such as GOT10k [54] and LaSOT [30], compared to baselines including Vision Mamba (linear), Agent Attention (sparse), and Mixformer [104] (quadratic). GOT10k LaSOT Model AO $\text{SR}_{0.5}$ $\text{SR}_{0.75}$ Suc. N-Pre. Pre. SA 0.704 0.796 0.675 0.690 0.785 0.749 Agent-A 0.695 0.787 0.662 0.644 0.731 0.689 mamba vision 0.700 0.789 0.673 0.677 0.771 0.730 Comba-$\operatorname{splr}$ 0.715 0.804 0.686 0.693 0.789 0.751 Comba-$\operatorname{iplr}$ 0.718 0.809 0.688 0.694 0.786 0.755
@@ 5 Conclusion & Future Work
Acknowledgments
Here, I would like to express my gratitude to those who helped me when I first entered this field, including my supervisor, my collaborators at Shanghai AI Lab, the members of the flash-linear-attention community, as well as Sonta and Zhangyu for their valuable assistance on this paper.
@@ A.1 Bilinear Systems
Formally, TTT, Gated-DeltaNet, RWKV-7, and Comba should be classified as bilinear systems [17, 113, 99, 102, 70]. These systems are linear with respect to both $\bm{S}$ and $\bm{K},\bm{V}$, but due to the interaction between $\bm{S}$ and $\bm{K}$, the overall system is nonlinear. Typically considered a special type of nonlinear system, bilinear systems possess more expressive power than linear systems while remaining more controllable than strictly nonlinear systems, such as chaotic systems. They are widely applied in the study of physical systems and biological population dynamics [102, 70].
@@ B.2 WY Representation & UT Transform
Detailed derivation can be found in the Appendix of DeltaNet [108, 105].
@@ B.3 Comba-SPLR-pk
Then, we employ the WY representation [13]:
@@
To maximize hardware efficiency, we apply the UT transform [51] to reduce non-matmul FLOPs, which is crucial to enable better hardware utilization during training.
@@
The inverse of a lower triangular matrix can be efficiently computed through an iterative row-wise approach by forward substitution in Gaussian elimination [37] and maintain data in $\operatorname{float32}$.

### v5

@@ 3.1 Forcing Forgetting for Long-range Modeling
Extensive work [lin2025forgetting, tan2024stick, press2021train] has shown that positional encoding is crucial for language models to generalize from short pretraining contexts to unconstrained inference lengths. While gated linear recurrent structures themselves can be viewed as the continuous accumulation of bias [tan2024stick], enabling these models to extrapolate effectively by learning how to forget. Recent studies [chen2024stuffed, ye2025longmamba, ben2024decimamba] suggest that pretraining lengths (e.g., 2K or 4K tokens) are insufficient to fill the model’s state capacity, thus making the forget gate close to 1, forcing the model to learn how to forget. As shown in Table LABEL:tab:forcing_forget, we evaluate representative gating initialization methods. For models like GLA [yang2023gated], the forget gate is initialized to $(0,1)$ and retains all incremental memories. For MetaLA [chou2024metala], HGRN2 [qin2024hgrn2], and GSA [zhang2024gated], the sum of the forget and input gates is constrained to 1, achieving a relative balance in system memory capacity. Experimental results in §4.2 indicate that breaking this balance is necessary. To this end, Comba follows Mamba2’s design by constraining the forget gate $\alpha_{t}$} to close to 1, while separately initializing the input gate $\beta_{t}$} to $(0,1)$ [yang2024parallelizing]. Additionally, we set the strength of state feedback correction is $\tilde{\beta}_{t}={b}\odot\beta_{t}$, where b{b} is computed by $\operatorname{Sigmoid}$ function to be constrained in interval $(0,1)$ to ensure that the feedback strength is weaker than the incremental information.
@@ 3.2 Comba with Chunk-wise Parallel
Based on Eq. 5, Comba can be implemented recursively, enabling constant memory usage and $\mathcal{O}(1)$ time complexity during the inference stage, where the Python-style pseudocode is provided in App. B.1. However, the naive recursive implementation in PyTorch lacks sufficient matrix multiplication and thread parallelism [dao2022flashattention], resulting in unacceptable overhead to pretraining. Referring to DeltaNet [yang2024gated], we optimize Comba through chunk-wise parallelism. App. B.3 presents an alternative form to fuse feedback decay factor $b$ into $\bm{k}$ as $\bm{p}=b\bm{k}$ for a flexible invocation.
@@
Eq. 6 involves matrix-matrix products at each time step, i.e., $\bm{S}_{[t]}^{0}@\bm{D}_{[t]}^{r}$, preventing parallelization in the sequence level. Then, we employ the WY representation [bischof1987wy] to eliminate these terms:
@@
To maximize hardware efficiency, we apply the UT transform [joffrain2006accumulating] to Eq. 7-8 to reduce non-matmul FLOPs, which is crucial to enable better hardware utilization during training:
@@
The inverse of a lower triangular matrix can be efficiently computed through an iterative row-wise approach by forward substitution in Gaussian elimination [grcar2011mathematicians] and maintain data in $\operatorname{float32}$. Notably, Comba computes the inverse matrix once in Eq. 10 (twice in the original Gated-DeltaNet). This is mainly attributed to the introduction of a new form of mathematical induction in deriving the WY representation in Eq. 7, leading to a more concise formulation and resulting in speedup [^3] During the forward pass, Comba alleviates the main bottleneck of inverse matrix computation with $\mathcal{O}(d^{3})$ in the original Gated-DeltaNet, yielding a 40% speedup. In the backward pass, performance gains diminish due to the reuse of cached inverse matrices $\bm{M}$. However, in large-scale models where recomputation is required, Comba is expected to offer significant performance advantages. Moreover, this structured modification applies to almost all models with Delta rule. Recently, inspired by Comba, Gated-DeltaNet has also been upgraded to a single inversion in flash-linear-attention, resulting in a significant speedup..
@@ 3.3 Neural Architecture
As shown in Fig. 2, we present the architecture for Comba families, where $\{a,b,c,d\}$ are trainable scalars (experimental results indicate that data dependency is not required). Following prior work [chou2024metala, yang2024gated, gu2023mamba, fu2022hungry], we introduce short convolutions to $\bm{qkv}$ to incorporate token shift to improve the model’s retrieval capacity. We also retain feature map operations [wang2020linformer] and utilize the $\operatorname{SiLU}$ function to approximate the exponential kernel in the softmax attention. To further stabilize training, we apply L2 normalization to $\bm{qk}$ and employ a $\operatorname{Sigmoid}$-based gating mechanism. Additionally, we explore a hybrid architecture [sun2024hunyuan, yang2024gated, ren2024samba] by integrating Comba layers directly with softmax attention.
@@ 4.1 Operator Efficiency Analysis
Figure 3: Operator speed evaluated on the Triton-Testing-Benchmark [tillet2019triton] (fwd and bwd) in single A800-80G GPU.
As shown in Fig. 3, we compared the speeds of various operators in both forward and backward processes. The recurrent Comba in PyTorch [paszke2017automatic] incurs significant computational overhead, limiting its scalability for large-scale pretraining. Flash-attention [dao2022flashattention] achieves the fastest speed for shorter sequences (e.g., 1024), but its quadratic complexity results in decreasing efficiency as sequence length increases. Among four modern RNN operators, Comba shows nearly 40% speed improvement in the forward process over Gated-DeltaNet due to a more efficient formula structure. GLA suffers from slower operator speed due to the use of diagonal gating matrices. Although RetNet achieves the fastest speed, it falls behind other models in performance due to the lack of data-dependent memory management (Table 6). Overall, Comba shows potential as a foundational framework.
@@ 4.2 Language Modeling & Architecture Ablation
Table 6: Zero-shot performance of 340M and 1.3B models trained on SlimPajama [cerebras2023slimpajama] datasets. The commonsense Reasoning task is evaluated by lm-evaluation-harness [eval-harness] and the recall-intensive task follows prefix-linear-attention [arora2024just] with 2K input tokens. * Some of the baseline results are from [yang2024parallelizing] and [du2025mom].
@@
As shown in the left half of Table 6, (i) most recursive models outperform transformers in commonsense reasoning tasks, due to their recursive structure that resembles a chain of thought [wei2022chain]. (ii) The SPLR structure outperforms both the IPLR and DPLR structures in two model sizes and achieves the highest computational efficiency. (iii) Output correction, i.e., $\left<q,dk\right$>, significantly reduces perplexity, enhancing memory utilization during question answering and improving performance across various metrics. (iv) We find that the Mamba architecture design is suboptimal. MLP, as a special (non-Hebbian) key-value memory [geva2020transformer, gershman2025key], complements the key-value associative memory in the state, which is particularly important for tasks such as inference. (v) Although Mamba’s multi-value attention model aids efficient key-value memory storage [li2024survey], it sacrifices performance compared to standard multi-head attention. (vi) We discover that the initialization of $d$ should be chosen differently for model scales, e.g., $d=0.02$ for 340M models and $d=1$ for 1.3B models, as smaller models are more prone to incorrect gradient descent directions.
@@
Inspired by previous hybrid architectures [yang2024gated, sun2024hunyuan, blakeman2025nemotron, dong2024hymba, li2025transmamba, chen2025transmamba], we empirically replace the second-to-last layer in every eight layers of Comba with softmax attention to boost the model’s recall ability. In Table 7, (i) we find that nearly all hybrid architectures outperform transformers even with a few quadratic layers. (ii) The best overall performance comes from the GSA-H [zhang2024gated], as GSA itself is a type of intra-layer hybrid architecture. Computationally, it can be expressed as two GLA operations followed by a softmax operation to address attention sparsity issues [tan2024stick], thus improving recall ability. In the future, we plan to combine GSA with Comba for a flexible hybrid.
@@
Table 8: Performance on LongBench [bai2024longbench] tasks with 10K length based on lm-evaluation-harness [eval-harness].
@@ 4.3 Vision Modeling
Table 9: Performance on the ImageNet-1K [deng2009imagenet] classification, compared to Vision Mamba [mambavision] (linear), DeiT [deit] (quadratic), and Agent Attention [agentattention] (sparse). Model Res. Params. FLOPs Top-1 DeiT-T 224*224 5.7(MB) 1.2(G) 72.2% Agent-T 224*224 6.0 1.2 74.9(+2.7) Vim-T 224*224 7.0 1.5 76.1(+3.9) Comba-T 224*224 5.8 1.1 76.3(+4.1) DeiT-S 224*224 22.1 4.6 79.8 Agent-S 224*224 22.7 4.6 80.5(+0.7) Vim-S 224*224 26.0 5.1 80.3(+0.5) Comba-S 224*224 22.6 4.4 80.5(+0.7)    Table 10: Performance on the object tracking datasets such as GOT10k [got10k] and LaSOT [lasot], compared to baselines including Vision Mamba (linear), Agent Attention (sparse), and Mixformer [mixformer] (quadratic). GOT10k LaSOT Model AO $\text{SR}_{0.5}$ $\text{SR}_{0.75}$ Suc. N-Pre. Pre. SA 0.704 0.796 0.675 0.690 0.785 0.749 Agent-A 0.695 0.787 0.662 0.644 0.731 0.689 mamba vision 0.700 0.789 0.673 0.677 0.771 0.730 Comba-$\operatorname{splr}$ 0.715 0.804 0.686 0.693 0.789 0.751 Comba-$\operatorname{iplr}$ 0.718 0.809 0.688 0.694 0.786 0.755
@@ 5 Conclusion & Future Work
Acknowledgments and Disclosure of Funding
Here, I would like to express my gratitude to those who helped me when I first entered this field, as well as the flash-linear-attention community for their valuable feedback on this paper.
This work is mainly supported by the Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515011994). This work is also supported by the National Natural Science Foundation of China (No. 62402414), Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Didi (CCF-DiDi GAIA Collaborative Research Funds), Guangzhou Municipal Science and Technology Project (No. 2023A03J0011), Huawei Industrial Funds, and the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628).
@@ A.1 Bilinear Systems
Formally, TTT, Gated-DeltaNet, RWKV-7, and Comba should be classified as bilinear systems [bruni1974bilinear, zhao2016gramian, wang2023expectation, williamson1977observation, pardalos2010optimization]. These systems are linear with respect to both $\bm{S}$ and $\bm{K},\bm{V}$, but due to the interaction between $\bm{S}$ and $\bm{K}$, the overall system is nonlinear. Typically considered a special type of nonlinear system, bilinear systems possess more expressive power than linear systems while remaining more controllable than strictly nonlinear systems, such as chaotic systems. They are widely applied in the study of physical systems and biological population dynamics [williamson1977observation, pardalos2010optimization].
@@ B.2 WY Representation & UT Transform
Detailed derivation can be found in the Appendix of DeltaNet [yang2024parallelizing, yang2024gated].
@@ B.3 Comba-SPLR-pk
Then, we employ the WY representation [bischof1987wy]:
@@
To maximize hardware efficiency, we apply the UT transform [joffrain2006accumulating] to reduce non-matmul FLOPs, which is crucial to enable better hardware utilization during training.
@@
The inverse of a lower triangular matrix can be efficiently computed through an iterative row-wise approach by forward substitution in Gaussian elimination [grcar2011mathematicians] and maintain data in $\operatorname{float32}$.